### v1

# Quantum $(r,\delta)$-Locally Recoverable Codes

Abstract.
A classical $(r,\delta)$-locally recoverable code is an error-correcting code such that, for each coordinate $c_{i}$ of a codeword, there exists a set of at most $r+\delta-1$ coordinates containing $c_{i}$ which allow us to correct any $\delta-1$ erasures in that set. These codes are useful for avoiding loss of information in large scale distributed and cloud storage systems.
In this paper, we introduce quantum $(r,\delta)$-locally recoverable codes and give a necessary and sufficient condition for a quantum stabilizer code $Q(C)$ to be $(r,\delta)$-locally recoverable. Our condition depends only on the puncturing and shortening at suitable sets of both the symplectic self-orthogonal code $C$ used for constructing $Q(C)$ and its symplectic dual $C^{\perp_{s}}$.
When a quantum stabilizer code comes from a Hermitian or Euclidean dual-containing code, and under an extra condition, we show that there is an equivalence between the classical and quantum concepts of $(r,\delta)$-local recoverability. A Singleton-like bound is stated in this case and examples of optimal stabilizer $(r,\delta)$-locally recoverable codes are given.
Key words and phrases:
Classical-quantum equivalence of local recovery, quantum stabilizer codes, Singleton-like bound
2020 Mathematics Subject Classification:
81P45; 94B65; 11T71
The first three authors were partially funded by MCICIU/AEI/10.13039/501100011033 and by “ERDF A way of making Europe” (grant PID2022-138906NB-C22), and by University Jaume I (grant GACUJIMA/2024/03). The third author was also partially funded by University Jaume I (grant PREDOC/2020/39). The fourth author was partially supported by “Japan Society of the Promotion of the Science” (grant 23K10980).
1. Introduction
Literature contains various sources giving some evidence of quantum supremacy [3, 6, 58, 36], which makes that quantum information processing have so much potential for multiple practical applications. When using quantum devices, it is important to protect quantum information from decoherence and other quantum noise. Quantum error-correcting codes (QECCs) are designed to perform this task since quantum correction works [49, 51] despite the fact that one cannot clone quantum information [16, 56].
Only binary codes were considered in the seminal contributions to the study of QECCs [11, 25, 12, 10]. At a later stage, $q$-ary QECCs ($q$ being a power of a prime number $p$) have been extensively studied and constructed with a wide variety of techniques, see for instance [8, 5, 29, 1, 28, 32, 19, 2] among many other papers on the subject.
A QECC of length $n$ is simply a $K$-dimensional linear subspace of the complex Hilbert space $\mathbb{C}^{q}\otimes\cdots\otimes\mathbb{C}^{q}=\mathbb{C}^{q^{n}}$. Quantum stabilizer codes constitute the most-known class of QECCs. Stabilizer codes are particularly interesting because of their relation with classical codes. In this paper, we use the definition of stabilizer code given in [5], where a commutative subgroup $S$ of the group $G_{n}$ generated by a nice error basis on $\mathbb{C}^{q^{n}}$ is considered. Then, the quantum stabilizer code defined by $S$ and a linear character $\lambda$ of $S$ is the set of common eigenvectors of elements $E\in S$ with eigenvalue $\lambda(E)$. The goodness of a quantum code is measured by its parameters $((n,K,d))_{q}$, where $d$ means that the code can detect all errors in $G_{n}$ whose weight is less than $d$ but not some of weight $d$.
With the above definition, we have the nice property that an $((n,K,d))_{q}$ stabilizer code exists if and only if there is an additive classical code $C\subseteq\mathbb{F}_{q}^{2n}$ which is self-orthogonal with respect to the trace-symplectic form. In addition, when $K>1$, $d$ coincides with the symplectic weight of $C^{\perp_{\mathrm{trs}}}\setminus C$, $C^{\perp_{\mathrm{trs}}}$ being the dual of $C$ with respect to the above form. This relation between quantum and classical error-correcting codes simplifies the obtention and study of stabilizer codes. If one uses the easier to handle Hermitian, or Euclidean, inner product on $\mathbb{F}_{q^{2}}^{n}$, or on $\mathbb{F}_{q}^{n}$, quantum stabilizer codes (and their parameters) can be gotten from self-orthogonal classical codes with respect to the mentioned products (and their parameters).
Some details on the definition and the way to get quantum stabilizer codes from classical codes are recalled in Subsection 2.2. The main results containing the parameters of the obtained codes are Theorems 6 and 7.
Many big companies store information in large scale distributed and cloud storage systems and, for them, the loss of the data in a node is a major issue. The so-called repair problem proposes to recover the lost data from the information deposited in the remaining nodes. To do it, in [24] the authors propose to use classical error-correcting codes (ECCs). Thus, we say that an ECC has $r$-locality when one can recover any coordinate of a word in $C$ by accessing at most $r$ other symbols of the word. These codes are named $r$-locally recoverable codes ($r$-LRCs) or LRCs with locality $r$. Some recent references on them are [53, 7, 34, 41, 38, 27, 35, 46, 13]. A Singleton-like bound for $r$-LRCs was proposed in [24], see (3). Optimal $r$-LRCs are those reaching that bound. Many of the above cited papers contain families of optimal LRCs.
Locally recoverable codes with locality $r$ are only designed for the case when a node fails, but simultaneous multiple device failures are not infrequent. To address this issue, $(r,\delta)$-locally recoverable codes ($(r,\delta)$-LRCs), or LRCs with locality $(r,\delta)$, were introduced in [44]. Fixed a coordinate $c_{i}$ of a codeword, these codes are capable of correcting $\delta-1$ erasures in a set containing $c_{i}$ plus, at most, $r+\delta-2$ other coordinates. These codes also admit a Singleton-like bound (4) and when reached they are called optimal $(r,\delta)$-LRCs. A considerable effort has been made recently to get optimal $(r,\delta)$-LRCs [15, 33, 38, 52, 27, 14, 17, 45, 9, 20]. Basic information on LRCs is given in Subsection 2.1.
The good prospects offered by quantum computing and the growing amount of information stored by data centers make it reasonable to introduce quantum locally recoverable codes. The large budget of companies such as Microsoft and Meta, could make it feasible to use quantum computers to store information, which would require local recovery. Quantum $r$-locally recoverable codes (quantum $r$-LRCs) were recently introduced in [23] for correcting an erasure corresponding to an input qudit. There, the authors give a condition for a stabilizer code to be a quantum $r$-LRC. Then, they particularize it to local recoverability of CSS codes. Despite the recent introduction of quantum $r$-LRCs, one can find contributions in this last line [37, 48]. In addition, [23], among other results, contains a Singleton-like bound, which admits a stronger version under a specific behavior of the recovery sets [23, Theorem 36]. Next, we reproduce that last bound for an $[[n,k,d]]_{q}$ quantum $r$-LRC:
(1) $k\leq\left(1-\frac{2}{r+1}\right)n-2\left(d-1-\left\lceil\frac{d-1}{r}\right\rceil\right).$
Given a QECC, $Q\subseteq\mathbb{C}^{q^{n}}$, and two sets of indices $I$ and $J$ such that $\emptyset\subsetneq I\subsetneq J\subseteq\{1,\ldots,n\}$, roughly speaking, we say that $Q$ is $(I,J)$-locally recoverable whenever $Q$ corrects erasures at (the positions in) $I$ from the qubits of $Q$ at $J$. See the more accurate Definition 8, where we introduce the concept of $(I,J)$-local recoverability with the aim of reverting the action of an arbitrary error channel (in fact, a completely depolarizing channel) on qubits at $I$ by using qudits at $J$.
In this paper, with the help of the previous definition, we introduce the concept of quantum $(r,\delta)$-local recoverability: A QECC $Q\subseteq\mathbb{C}^{q^{n}}$ is a quantum $(r,\delta)$-locally recoverable code (quantum $(r,\delta)$-LRC) if for each index $i\in\{1,\ldots,n\}$, there exists a set $J\subseteq\{1,\ldots,n\}$ containing $i$ with $\mathrm{card}\,(J)\leq r+\delta-1$ such that, for all subset $I\subseteq J$ of cardinality $\delta-1$, $Q$ is $(I,J)$-locally recoverable.
In Section 3, we study quantum stabilizer $(r,\delta)$-LRCs, $Q(C)$. That is, we consider stabilizer codes given by symplectic self-orthogonal codes $C\subseteq\mathbb{F}_{q}^{2n}$; that is $C\subseteq C^{\perp_{s}}$, where duality is with respect to the symplectic form defined in (7). Considering the concepts of shortening of a classical code $D$, at the set $I$, denoted $\sigma_{I}(D)$, and puncturing of $D$ at $J$, denoted $\pi_{J}(D)$, Theorem 17 gives a necessary and sufficient condition on the code $C$ and its symplectic dual for $Q(C)$ to be $(I,J)$-locally recoverable. Note that, contrary to usual, $I$ and $J$ refers to the coordinates of the codewords we keep (see the beginning of Section 2). Theorem 17 allows us to prove Propositions 19 and 20 which supply conditions implying or discarding $(I,J)$-local recoverability of $Q(C)$. These conditions depend only on the cardinality of $I$ and $J$ and certain generalized symplectic weights of the symplectic dual of $C$.
As a consequence of Theorem 17, we deduce the main result in the paper:
Theorem A (Theorem 21 in the paper).
Let $Q(C)\subseteq\mathbb{C}^{q^{n}}$ be a stabilizer code given by a symplectic self-orthogonal $q$-ary linear code $C\subseteq\mathbb{F}_{q}^{2n}$, $q$ being a prime power. $Q(C)$ is a quantum $(r,\delta)$-LRC if and only if, for all index $i\in\{1,\ldots,n\}$, there exists a set $J\ni i$ of cardinality $\leq r+\delta-1$ such that, for any subset $\emptyset\neq I\subsetneq J$ of cardinality $\delta-1$, the following equality holds:
$\sigma_{I}\left[\pi_{J}\left(C^{\perp_{s}}\right)\right]=\sigma_{I}(C).$
Section 4 is devoted to specialize the results in Section 3 to QECCs coming from Hermitian or Euclidean self-orthogonal codes $C$ included in $\mathbb{F}_{q^{2}}^{n}$ or $\mathbb{F}_{q}^{n}$, respectively. As can be expected, a set of equalities as in Theorem A, where one replaces symplectic dual with Hermitian (respectively, Euclidean) dual, give a necessary and sufficient condition for $Q(C)$ being an $(r,\delta)$-LRC.
Replacing self-orthogonality of the involved classical codes with dual-containing, our second main result determines a condition under which stabilizer $(r,\delta)$-LRCs and classical $(r,\delta)$-LRCs are equivalent objects. Let us state the result, where we stand $Q^{\prime}(C)$ for the quantum code coming from a Hermitian or Euclidean dual-containing classical code $C$.
Theorem B (Theorem 28 in the paper).
Let $C$ be a $q^{2}$-ary linear code included in $\mathbb{F}_{q^{2}}^{n}$ (respectively, a $q$-ary linear code included in $\mathbb{F}_{q}^{n}$). Assume that $C$ is Hermitian (respectively, Euclidean) dual-containing. Suppose also that $\delta\leq d(C^{\perp_{h}})$ (respectively, $\delta\leq d(C^{\perp_{e}})$). Then, $Q^{\prime}(C)$ is a quantum stabilizer $(r,\delta)$-LRC if and only if $C$ is a classical $(r,\delta)$-LRC.
Then, under the condition $\delta\leq d(C^{\perp_{h}})$ (respectively, $\delta\leq d(C^{\perp_{e}})$), stated in Theorem B, we propose in Theorem 30 the following Singleton-like bound for quantum $(r,\delta)$-LRCs, $Q^{\prime}(C)$, with parameters $[[n,k,\geq d(C)]]_{q}$:
(2) $k+2d(C)+2\left(\left\lceil\frac{n+k}{2r}\right\rceil-1\right)(\delta-1)\leq n+2.$
A quantum code $Q^{\prime}(C)$, as above, is named to be optimal if it attains the above bound. Making use of Theorem B, Section 5 contains some results implying optimality of certain families of stabilizer $(r,\delta)$-LRCs. In Subsection 5.2, we show that Hermitian dual-containing $q^{2}$-ary MDS codes give rise to optimal stabilizer $(r,\delta)$-LRCs and that the same happens with the Euclidean case. Other families of optimal stabilizer $(r,\delta)$-LRCs coming from $\emptyset$-affine variety codes are also given in Subsection 5.1. Finally, setting $\delta=2$, we find an optimal quantum $(r,2)$-LRC which, regarded as a quantum $r$-LRC, attains the bound (1).
2. Preliminaries
In this paper, we consider error-correcting linear codes $C\subseteq\mathbb{F}_{q}^{n}$ over a finite field of cardinality $q=p^{m}$, $p$ being a prime number a $m$ a positive integer. We also assume that the parameters of $C$ are $[n,k,d]_{q}$. We recall the concepts of shortening and puncturing of a code at a set $R$; note that in our definitions, $R$ refers to the coordinates of the codewords we keep.
Set $R\subseteq\{1,\ldots,n\}$, assume that $r=\mathrm{card}\,R$ and consider the projection map on the coordinates of $R$, $\pi_{R}:\mathbb{F}_{q}^{n}\rightarrow\mathbb{F}_{q}^{r}$. We define the puncturing of $C$ at $R$ as the set
$\pi_{R}(C)=\{\mathbf{c}_{R}:=(c_{j})_{j\in R}\;:\;\mathbf{c}=(c_{1},\ldots,c_{n})\in C\},$
$\pi_{R}(C)$ is also named the punctured code of $C$ at $R$. We also define the shortening of $C$ at $R$ as the set
$\sigma_{R}(C)=\{\mathbf{c}_{R}\;:\;\mathbf{c}=(c_{1},\ldots,c_{n})\in C\mbox{ satisfying supp}(\mathbf{c})\subseteq R\},$
where supp$(\mathbf{c}):=\{j\in\{1,\ldots,n\}:c_{j}\neq 0\}$. We also say that $\sigma_{R}(C)$ is the shortened code of $C$ at $R$. By [43], it holds that
$\pi_{R}\left(C^{\perp_{e}}\right)=\left[\sigma_{R}\left(C\right)\right]^{\perp_{e}}.$
We also remind that, for each index $1\leq t\leq k$, the value
$\omega_{t}(C):=\min\{\mathrm{card}\,[\mathrm{supp}(C^{\prime})]:C^{\prime}\mbox{ is a $t$-dimensional subcode of }C\}$
is named the $t$th generalized Hamming weight of $C$ [55, 26].
Our first subsection introduces locally recoverable codes and some known facts on them.
2.1. Classical $(r,\delta)$-locally recoverable codes
Data loss in large scale distributed and cloud storage systems is an important problem for many companies like Google, Meta, Microsoft, etc. In many cases, the problem occurs due to failures in one of its nodes. Locally recoverable codes (LRCs) were introduced in [24] to minimize this problem and protect the data. LRCs are classical error-correcting codes allowing one or more erased coordinates in a codeword to be recovered from other coordinates in the codeword.
We start with the classical definition of LRC. Let $C\subseteq\mathbb{F}_{q}^{n}$ be a linear $q$-ary code as before and set $i\in\{1,\ldots,n\}$ and $R\subseteq\{1,\ldots,n\}$ such that $i\notin R$. $R$ is said to be a recovery set for the $i$th coordinate of $C$ whenever an erasure in the coordinate $c_{i}$ of a codeword $\mathbf{c}=(c_{1},\ldots,c_{n})\in C$ can be recovered from the coordinates of $\mathbf{c}$ with indices in $R$. This allows us to define the locality of the $i$th coordinate in $C$ as the smallest cardinality of a recovery set for the $i$th coordinate and give the following definition:
Definition 1.
An error-correcting code $C$ is an $r$-LRC (or an LRC with locality $r$) whenever any coordinate of $C$ admits a recovery set, $r$ being the largest locality of its coordinates.
The following twofold result is well-known. See [20, Proposition 2.1] and [22, Proposition 1.1].
Proposition 2.
With the above notation the following results hold.
  1. (1)
    A set of coordinates $R\subseteq\{1,\ldots,n\}$ is locally recoverable for the $i$th coordinate of a code $C$ of length $n$, $i\notin R$, if and only if the minimum (Hamming) distance $d$ of the code $\pi_{\bar{R}}(C)$, where $\bar{R}:=R\cup\{i\}$, satisfies $d[\pi_{\bar{R}}(C)]\geq 2$.
  2. (2)
    Let $r$ be the locality of an LRC, $C$, then $r\geq d(C^{\perp_{e}})-1$, $C^{\perp_{e}}$ being the Euclidean dual of $C$.
As shown in [24], parameters $[n,k,d]_{q}$ and locality $r$ of an LRC code satisfy a Singleton-like bound:
(3) $k+d+\left\lceil{\frac{k}{r}}\right\rceil\leq n+2.$
Optimal LRCs are those meeting the bound in the above inequality.
As shown in Proposition 2, LRCs correct only an erasure and thus, they are only useful when a node of the system fails. However simultaneous failures are not unusual. The so-called $(r,\delta)$-LRCs are designed to address this problem. They were introduced for the first time in [44]. Let us define this class of LRCs.
Definition 3.
An error-correcting code $C\subseteq\mathbb{F}_{q}^{n}$ is an $(r,\delta)$-LRC (or an LRC with locality $(r,\delta)$) if for each index $i\in\{1,\ldots,n\}$, there exists a set of indices $J\subseteq\{1,\ldots,n\}$ that contains $i$ and satisfies the following two conditions:
  1. (1)
    $\mathrm{card}\,(J)\leq r+\delta-1$.
  2. (2)
    $d[\pi_{J}(C)]\geq\delta$.
The set $J$ is named an $(r,\delta)$-recovery set for the $i$th coordinate of $C$. Our definition implies that, for any set $I\subsetneq J$ of cardinality $\delta-1$, the coordinates corresponding to $I$ can be corrected from the knowledge of the coordinates corresponding to $J\setminus I$.
A Singleton-like bound has also been introduced for $(r,\delta)$-LRCs [44]. Let us state it.
Proposition 4.
Let $C$ be an $(r,\delta)$-LRC with parameters $[n,k,d]_{q}$. Then, the following inequality holds.
(4) $k+d+\left(\left\lceil{\frac{k}{r}}\right\rceil-1\right)\left(\delta-1\right)\leq n+1.$
As before, an optimal $(r,\delta)$-LRC is an $(r,\delta)$-LRC that meets the bound in Proposition 4. In this paper, we are interested in $(r,\delta)$-LRCs, therefore, unless otherwise specified, the sentence optimal (classical) code will stand for optimal $(r,\delta)$-LRC. In recent years there has been great activity in the search of optimal LRCs. Some interesting references are [50, 57, 31, 9], where one can find tables describing the known parameters on its publishing date of optimal codes.
We conclude this subsection by noting that, as stated in [22, Corollary 1.5], the above introduced concept of generalized Hamming weight allows us to give a result extending that stated in Part (2) of Proposition 2:
Proposition 5.
Let $C$ be an $(r,\delta)$-LRC. Then $r+\delta\geq\omega_{\delta-1}(C^{\perp_{e}})+1$.
Our last subsection briefly recalls basic information on quantum stabilizer codes.
2.2. Quantum stabilizer codes
Quantum information has to be protected from decoherence and other quantum noise, and quantum error-correcting codes (QECCs) are the main instrument to carry it out. We are interested in the study of quantum stabilizer locally recoverable codes. To introduce quantum stabilizer codes, we keep the above notation and use the definition and ideas given in [5].
A linear character of a finite commutative group $S$ is a group homomorphism from $S$ to $\mathbb{C}^{*}$, where $\mathbb{C}^{*}$ denotes the multiplicative group of the complex numbers. Set $\xi=e^{\frac{\iota 2\pi}{p}}$, $\iota=\sqrt{-1}$, a primitive $p$th root of unity, and consider a nice error basis on $\mathbb{C}^{q^{n}}$, $\epsilon_{n}:=\{E_{(\mathbf{a},\mathbf{b})}\;:\;(\mathbf{a}|\mathbf{b})\in\mathbb{F}_{q}^{2n}\}$, where $E_{(\mathbf{a},\mathbf{b})}$ are error operators obtained by tensoring generalized Pauli matrices acting on $\mathbb{C}^{q}$, as defined in [5]. Consider the group
$G_{n}:=\{\xi^{\ell}E_{(\mathbf{a},\mathbf{b})}\;:\;(\mathbf{a}|\mathbf{b})\in\mathbb{F}_{q}^{2n},0\leq\ell\leq p-1\}.$
[…]
$Q=Q(S):=\bigcap_{E\in S}\{\mathbf{v}\in\mathbb{C}^{q^{n}}\;:\;E\mathbf{v}=\lambda(E)\mathbf{v}\}.$
Any code $Q$ as before is associated to a character but, unless necessary, we will not mention it.
Setting $\mathbf{a}=(a_{1},\ldots,a_{n})$ and $\mathbf{b}=(b_{1},\ldots,b_{n})$, $a_{j},b_{j}\in\mathbb{F}_{q}$, the symplectic weight of an element $(\mathbf{a}|\mathbf{b})\in\mathbb{F}_{q}^{2n}$ is defined as
$\mathrm{swt}(\mathbf{a}|\mathbf{b}):=\mathrm{card}\,\{j:(a_{j},b_{j})\neq(0,0)\}.$
In addition, the weight $\mathrm{wt}(E)$ of $E=\xi^{\ell}E_{(\mathbf{a},\mathbf{b})}$ is defined as $\mathrm{swt}(\mathbf{a}|\mathbf{b})$. Finally, $Q$ has minimum distance $d$ if it detects all errors in $G_{n}$ of weight less than $d$ but not some of weight $d$. An $((n,K,d))_{q}$ quantum stabilizer code is a $K$-dimensional subspace of $\mathbb{C}^{q^{n}}=\mathbb{C}^{q}\otimes\cdots\otimes\mathbb{C}^{q}$ with minimum distance $d$. In case $K=q^{k}$, it is said to be an $[[n,k,d]]_{q}$ stabilizer code.
A quantum stabilizer code defines a linear code of length $2n$
$C:=\{(\mathbf{a}|\varphi^{-1}\mathbf{b})\;:\;\xi^{\ell}E_{(\mathbf{a},\mathbf{b})}\in S\},$
for a convenient automorphism $\varphi$ of $\mathbb{F}_{q}$ which is translated componentwise to $\mathbb{F}_{q}^{n}$. The automorphism $\varphi$ is chosen so that $C$ is self-orthogonal with respect to the following trace-symplectic form:
$(\mathbf{a}|\mathbf{b})._{\mathrm{trs}}(\mathbf{c}|\mathbf{d})=\mathrm{tr}_{q/p}(\mathbf{a}._{e}\mathbf{d}-\mathbf{b}._{e}\mathbf{c}),$
and we usually omit it.
Given a $q$-ary additive code $C\subseteq\mathbb{F}_{q}^{n}$, we denote by $C^{\perp_{\mathrm{trs}}}$ the dual code of $C$ with respect to the trace-symplectic form. The main theorem in this regard is the following one:
Theorem 6.
The existence of an $((n,K,d))_{q}$ stabilizer code is equivalent to that of an additive code $C\subseteq\mathbb{F}_{q}^{2n}$ with cardinality $q^{n}/K$, which is self-orthogonal with respect to the trace-symplectic form, and, whenever $K>1$, $d$ is the minimum symplectic weight of the elements in $C^{\perp_{\mathrm{trs}}}\setminus C$. When $K=1$, there is no undetectable error.
This result has two useful consequences that we state next. Firstly, recall that for $\mathbf{a},\mathbf{b}\in\mathbb{F}_{q^{2}}^{n}$, the Hermitian inner product is defined as $\mathbf{a}\cdot_{h}\mathbf{b}:=\sum_{j=1}^{n}a_{j}b_{j}^{q}$. Given a $q^{2}$-ary linear code $C$, $C^{\perp_{h}}$ denotes its dual with respect to $\cdot_{h}$. We stand $\mathrm{wt}$ for Hamming weight.
Theorem 7.
The following two items hold:
  1. (1)
    The existence of an $[n,k,d]_{q^{2}}$ linear code $C$ which is Hermitian dual-containing (that is $C^{\perp_{h}}\subseteq C$) implies that of an $[[n,2k-n,\geq d]]_{q}$ quantum stabilizer code.
  2. (2)
    The existence of two $q$-ary linear codes $C_{j}$ with parameters $[n,k_{j},d_{j}]_{q}$, $1\leq j\leq 2$, such that $C_{2}^{\perp_{e}}\subseteq C_{1}$ implies the existence of an $[[n,k_{1}+k_{2}-n,d]]_{q}$ quantum stabilizer code with $d=\min\{\mathrm{wt}(\mathbf{z}):\mathbf{z}\in(C_{1}\setminus C_{2}^{\perp_{e}})\cup(C_{2}\setminus C_{1}^{\perp_{e}})\}$.
When stating our results we have followed the terminology in [29], however our definition of stabilizer codes comes from [5]. There, the results are initially stated and proved over $\mathbb{F}_{p}$ and later extended to $\mathbb{F}_{q}$. Details of how to perform that extension can be found in [21, Section 3.1] in the context of entanglement-assisted codes. In this paper we will use this procedure.
We conclude this subsection by explaining when a set of erasures is correctable by a quantum code $Q$. Stand $A^{\dagger}$ for the conjugate transpose (or Hermitian conjugate) of a complex matrix $A$. As before, denote by $\epsilon_{1}$ the set of (generalized) Pauli matrices acting on $\mathbb{C}^{q}$ and, for a $q\times q$ complex matrix $\rho$, consider the completely depolarizing channel on a single qudit in $\mathbb{C}^{q}$:
$\Gamma(\rho)=\frac{1}{q^{2}}\sum_{E\in\epsilon_{1}}E\rho E^{\dagger}.$
Note that, with $\Gamma$, every possible error in $\epsilon_{1}$ (including the $q\times q$ identity matrix $\mathcal{I}_{q}$) is applied with probability $1/q^{2}$. Consider now a quantum code $Q\subseteq\mathbb{C}^{q^{n}}=\mathbb{C}^{q}\otimes\cdots\otimes\mathbb{C}^{q}$ and a set of indices $I\subseteq\{1,\ldots,n\}$. Set $\Gamma^{I}$ the map $\bigotimes_{j=1}^{n}\Gamma_{j}$ where $\Gamma_{j}=\Gamma$ when $j\in I$ and $\Gamma_{j}=\mathcal{I}_{q}$ otherwise; that is the map applying $\Gamma$ on all qudits at (the positions in) $I$. Then, by definition, erasures at $I$ are correctable by $Q$ if there exists a recovery operator $\mathcal{R}_{Q,I}$, which is a trace-preserving quantum operation (in the sense of [42, Chapter 8]), such that
(5) $\mathcal{R}_{Q,I}\circ\Gamma^{I}(|\varphi\rangle\langle\varphi|)=|\varphi\rangle\langle\varphi|$
for all $|\varphi\rangle\in Q$.
3. Quantum $(r,\delta)$-locally recoverable codes
3.1. Definitions
We start this section by introducing the concept of quantum code $Q\subseteq\mathbb{C}^{q^{n}}$ with alphabet size $q$ that corrects erasures at the positions in $I\subseteq\{1,\ldots,n\}$ from the qubits at a set $J\supsetneq I$, $J\subseteq\{1,\ldots,n\}$. Recall the definition of $\Gamma^{I}$ given above.
Definition 8.
Given two sets of indices as before $\emptyset\neq I\subsetneq J\subseteq\{1,\ldots,n\}$, a quantum code $Q\subseteq\mathbb{C}^{q^{n}}$ is said to be an $(I,J)$-locally recoverable code if there exists a trace-preserving quantum operation $\mathcal{R}_{Q,I}^{J}$, which acts only on the qubits corresponding to $J$ and keeps untouched the remaining ones, such that
(6) $\mathcal{R}_{Q,I}^{J}\circ\Gamma^{I}(|\varphi\rangle\langle\varphi|)=|\varphi\rangle\langle\varphi|$
for all $|\varphi\rangle\in Q$.
This definition allows us to introduce the concept of quantum $(r,\delta)$-locally recoverable code.
Definition 9.
A quantum code $Q\subseteq\mathbb{C}^{q^{n}}$ is said to be a quantum $(r,\delta)$-locally recoverable code (quantum $(r,\delta)$-LRC) if for each index $i\in\{1,\ldots,n\}$, there exists a set $J\subseteq\{1,\ldots,n\}$ containing $i$ with $\mathrm{card}\,(J)\leq r+\delta-1$ satisfying that for all subset $I\subseteq J$ of cardinality $\delta-1$, $Q$ is $(I,J)$-locally recoverable.
3.2. $(I,J)$-local recoverability for quantum stabilizer codes
In this paper, for constructing stabilizer codes, we use a slightly more restrictive version of the trace-symplectic form. Indeed, for $(\mathbf{a}|\mathbf{b}),(\mathbf{c}|\mathbf{d})\in\mathbb{F}_{q}^{2n}$, we consider the following symplectic form:
(7) $(\mathbf{a}|\mathbf{b})\cdot_{s}(\mathbf{c}|\mathbf{d}):=\mathbf{a}\cdot_{e}\mathbf{d}-\mathbf{b}\cdot_{e}\mathbf{c}.$
From now on, $Q(C)$ denotes the quantum stabilizer code defined by a symplectic self-orthogonal (with respect to $\cdot_{s}$) classical code $C\subseteq\mathbb{F}_{q}^{2n}$; that is $C\subseteq C^{\perp_{s}}$, where the duality is with respect the symplectic form $\cdot_{s}$.
We start by explaining a standard procedure for correcting erasures at $I\subsetneq\{1,\ldots,n\}$ of the code $Q(C)$. Our procedure is the same for codes $C\subseteq\mathbb{F}_{q^{2}}^{n}$ (respectively, $C\subseteq\mathbb{F}_{q}^{n}$) such that $C\subseteq C^{\perp_{h}}$ (respectively, $C\subseteq C^{\perp_{e}}$).
Let $[[n,k]]_{q}$ be the parameters of $Q(C)$, then $\dim C=n-k$ and $\dim(C^{\perp_{s}})=n+k$. Consider a basis $\{\mathbf{h}_{1},\ldots,\mathbf{h}_{n-k}\}$ of the linear space $C$ and the surjective map
$f:\mathbb{F}_{q}^{2n}\rightarrow\mathbb{F}_{q}^{n-k};\;\;f(\mathbf{y})=(\mathbf{y}\cdot_{s}\mathbf{h}_{1},\ldots,\mathbf{y}\cdot_{s}\mathbf{h}_{n-k})$
whose kernel is $C^{\perp_{s}}$. Then, in a first step, for each index $1\leq\ell\leq n-k$, the decoder measures an observable whose eigenspaces are the same as $E_{\mathbf{h}_{\ell}}$ and the measurement outcomes provide the vector $(s_{1},\ldots,s_{n-k})\in\mathbb{F}_{q}^{n-k}$, where $s_{\ell}=\mathbf{e}\cdot_{s}\mathbf{h}_{\ell}$, for some error vector $\mathbf{e}\in\mathbb{F}_{q}^{2n}$. In a second step, setting $\mathbf{y}=(\mathbf{a},\mathbf{b})$ an unknown, the decoder considers the linear system of equations
(8) $f(\mathbf{y})=(s_{1},\ldots,s_{n-k});\;a_{j}=b_{j}=0,j\not\in I.$
Finally, since we desire to correct erasures at $I$, if the solution $\mathbf{e}$ of the system is unique modulo $C$, then the decoder applies the inverse $E_{\mathbf{e}}^{-1}$. Otherwise, the decoder declares failure.
We take into account the above procedure, to state the following result.
Theorem 10.
Let $C$ be a symplectic self-orthogonal classical $\mathbb{F}_{q}$-linear code and $Q(C)$ the corresponding stabilizer code. Keep also the above notation. Then, the equality
(9) $\sigma_{I}(C)=\sigma_{I}(C^{\perp_{s}})$
is a necessary and sufficient condition for $Q(C)$ to correct erasures at $I$.
Proof.
By [39, Equation (3)], the system of linear equations (8) has a unique solution when Equality (9) holds, and therefore (9) implies that $Q(C)$ corrects erasures at $I$.
To conclude the proof, we are going to show that if Equality (9) is false, then no decoding procedure can correct erasures at $I$. Indeed, if Equality (9) does not hold, there is an erasure $\mathbf{e}=(\mathbf{a},\mathbf{b})\in C^{\perp_{s}}\setminus C$ with $a_{j}=b_{j}=0$ for $j\not\in I$. Then $E_{\mathbf{e}}$ sends some quantum codeword in $Q(C)$ to another one in $Q(C)$, which prevents to correct all erasures at $I$ by any decoding procedure (in the sense of (5)). ∎
In the rest of this subsection we study the case $J\subsetneq\{1,\ldots,n\}$. We divide our study in two parts. The first one provides a sufficient condition for a stabilizer code to be $(I,J)$-locally recoverable. In the second part, we will show that the condition stated in the first part is also necessary.
As before mentioned, we state and prove our results for the case when the alphabet size is a prime number $p$; this is because they can be extended to a general $q$ following the reasoning given in [21, Section 2.1].
3.2.1. A sufficient condition
We desire to correct erasures from qudits at a set $J\subsetneq\{1,\ldots,n\}$, therefore we are interested in what happens when considering codewords in $Q(C)$ and discarding their qudits at $\overline{J}=\{1,\ldots,n\}\setminus J$. Thus let $|\varphi\rangle\in Q(C)$ be a codeword in our stabilizer code, the partial trace over $\overline{J}$, $\sigma=\mathrm{tr}_{\overline{J}}[|\varphi\rangle\langle\varphi|]$, of the density matrix $|\varphi\rangle\langle\varphi|$ allows us to consider the part of the codeword $|\varphi\rangle$ in $J$.
In our case, where $C\subseteq\mathbb{F}_{p}^{2n}$, for a vector $\mathbf{y}=(\mathbf{a}|\mathbf{b})$ we denote $\pi_{J}(\mathbf{y})=(a_{j}|b_{j})_{j\in J}$ and, as before, the shortening of $C$ at $J$ is defined as
$\sigma_{J}(C)=\{\pi_{J}(\mathbf{y})\;:\;\mathbf{y}\in C\mbox{ and }\mathrm{supp}(\mathbf{a}),\mathrm{supp}(\mathbf{b})\subseteq J\}.$
Analogously, the puncturing of $C$ at $J$ is $\pi_{J}(C)=\{\pi_{J}(\mathbf{y})\;:\;\mathbf{y}\in C\}$. Then, it is clear that $|\varphi\rangle$ is an eigenvector of an observable $E_{\mathbf{y}}$, where $\mathbf{y}$ is given by a vector $\mathbf{z}\in\sigma_{J}(C)$ to which we add zeroes at the coordinates supported in $\overline{J}$. Our first result relates the column space of the above density matrix, $\mathrm{col}(\sigma)$, to the stabilizer code defined by the shortened code $\sigma_{J}(C)$.
Proposition 11.
Let $Q(C)\subseteq\mathbb{C}^{p^{n}}$ be the quantum stabilizer code defined by a symplectic self-orthogonal classical code $C$. Let $J\subsetneq\{1,\ldots,n\}$ be a set of indices. Consider an element $|\varphi\rangle\in Q(C)$ and the density matrix $\sigma=\mathrm{tr}_{\overline{J}}[|\varphi\rangle\langle\varphi|]$. Set $Q(\sigma_{J}(C))\subseteq\mathbb{C}^{p^{\mathrm{card}(J)}}$ the stabilizer code defined by the shortened code $\sigma_{J}(C)$ such that, for each pair of codewords $\mathbf{y}\in C$ and $\mathbf{z}\in\sigma_{J}(C)$ behaving as indicated before the statement, with respective observables $E_{\mathbf{y}}$ and $E_{\mathbf{z}}$, the characters $\lambda(E_{\mathbf{y}})$ and $\lambda(E_{\mathbf{z}})$ coincide. Then, the column space, $\mathrm{col}(\sigma)$, satisfies
$\mathrm{col}(\sigma)\subseteq Q(\sigma_{J}(C)).$
Proof.
It is clear that $\sigma_{J}(C)\subseteq\mathbb{F}_{p}^{2\,\mathrm{card}(J)}$. Denote by $P(\mathbf{y})$ the projection matrix of $E_{\mathbf{y}}$ on $\mathbb{C}^{p^{n}}$ corresponding to the measurement outcome $\lambda(E_{\mathbf{y}})$. Then, the probability of not getting $\lambda(E_{\mathbf{y}})$ equals
$1-\mathrm{tr}[|\varphi\rangle\langle\varphi|P(\mathbf{y})]=1-\mathrm{tr}[\sigma P(\mathbf{z})]=\mathrm{tr}[\sigma(\mathcal{I}_{p^{\mathrm{card}(J)}}-P(\mathbf{z}))]=0,$
where $\mathrm{tr}$ stands for the trace map and $\mathcal{I}_{p^{\mathrm{card}(J)}}$ is the identity matrix of size $p^{\mathrm{card(J)}}\times p^{\mathrm{card(J)}}$. As a consequence, and since $\sigma$ is positive semidefinite, one gets that $\mathrm{col}(\sigma)\subseteq\mathrm{col}(P(\mathbf{z}))$.
The relation between $\mathbf{z}$ and $\mathbf{y}$ and the fact that $Q(\sigma_{J}(C))$ is the intersection of column spaces $\mathrm{col}(E_{\mathbf{z}})$, where $\mathbf{z}$ runs over $\sigma_{J}(C)$, concludes the proof. ∎
Let us state our main result in this subsection.
Proposition 12.
Let $Q(C)\subseteq\mathbb{C}^{p^{n}}$ be a stabilizer code as above and consider sets of indices $\emptyset\neq I\subsetneq J\subsetneq\{1,\ldots,n\}$. The codes’ equality
(10) $\sigma_{I}\left[\pi_{J}\left(C^{\perp_{s}}\right)\right]=\sigma_{I}(C)$
is a sufficient condition for $Q(C)$ to be $(I,J)$-locally recoverable.
Proof.
Consider a codeword $|\varphi\rangle\in Q(C)$ for which qudits at $I$ were erased. We desire to recover $|\varphi\rangle$ (in fact, the qubits at $I$) by measurements and a unitary operator on qudits at $J$. By Proposition 11, every vector in $\mathrm{col}(\sigma)$, $\sigma=\mathrm{tr}_{\overline{J}}[|\varphi\rangle\langle\varphi|]$, is a codeword in $Q(\sigma_{J}(C))$. Then, to recover erasures at $I$, we can use the above defined decoder $\mathcal{R}_{Q,I}$ (for the quantum code $Q=Q(\sigma_{J}(C))$). By Theorem 10, $Q(\sigma_{J}(C))$ correct erasures at $I$ if and only if
$\sigma_{I}\left[\left(\sigma_{J}(C)\right)^{\perp_{s}}\right]=\sigma_{I}\left[\sigma_{J}(C)\right],$
which is equivalent to
$\sigma_{I}\left[\left(\sigma_{J}(C)\right)^{\perp_{s}}\right]=\sigma_{I}(C),$
and, by [43], this last equality is equivalent to
$\sigma_{I}\left[\pi_{J}\left(C^{\perp_{s}}\right)\right]=\sigma_{I}(C).$
This concludes the proof. ∎
3.2.2. A necessary and sufficient condition for $(I,J)$-local recoverability
The aim of this section is to show that Equality (10) is not only a sufficient but also a necessary condition for $Q(C)$ to be an $(I,J)$-locally recoverable code. Our proof will be a consequence of the forthcoming Lemmas 13, 14 and 15.
Our first result shows the necessity of Equality (10) as long as certain requirement holds. Before to state it, we recall that a completely mixed state on an $\ell$-dimensional linear space $V\subseteq\mathbb{C}^{p^{n}}$ is an equal probabilistic mixture of the elements of an orthonormal basis of $V$.
Lemma 13.
Let $Q(C)\subseteq\mathbb{C}^{p^{n}}$ be a stabilizer code given by a symplectic self-orthogonal $p$-ary code $C\subseteq\mathbb{F}_{p}^{2n}$, and consider sets of indices $\emptyset\neq I\subsetneq J\subsetneq\{1,\ldots,n\}$. Set $\overline{J}=\{1,\ldots,n\}\setminus J$. Suppose that there exists an environment $\mathcal{E}$ surrounding $\mathbb{C}^{p^{n}}$ and an entangled quantum codeword $|\Phi\rangle\in Q(C)\otimes\mathcal{E}$ such that the partial trace of $|\Phi\rangle\langle\Phi|$ over $\overline{J}$ and $\mathcal{E}$ is a completely mixed state on $Q(\sigma_{J}(C))$. Then the codes’ equality stated in (10):
$\sigma_{I}\left[\pi_{J}\left(C^{\perp_{s}}\right)\right]=\sigma_{I}(C)$
is a necessary condition for $Q(C)$ to be $(I,J)$-locally recoverable.
Proof.
We reason by contradiction and thus we suppose that Equality (10) is false. Then, there exists some element $\mathbf{e}\in\pi_{J}(C^{\perp_{s}})$ with zero coordinates out of $I$ such that $\pi_{I}(\mathbf{e})\notin\sigma_{I}(C)$. Then $E_{\mathbf{e}}|\xi\rangle\in Q(\sigma_{J}(C))$ for all $|\xi\rangle\in Q(\sigma_{J}(C))$. Since $\mathbf{e}$ is not in the stabilizer of $Q(\sigma_{J}(C))$, there exists $|\zeta\rangle\in Q(\sigma_{J}(C))$ such that $E_{\mathbf{e}}|\zeta\rangle$ is not a scalar multiple of $|\zeta\rangle$ and it is physically different from $|\zeta\rangle$.
Suppose that a local recovery procedure $\mathcal{S}:=\mathcal{S}^{J}_{Q(\sigma_{J}(C)),I}$ keeps $E_{\mathbf{e}}|\zeta\rangle$ when no erasure happens. Then $\mathcal{S}$ does not correct $|\zeta\rangle$ when the erasure $\mathbf{e}$ happens as $\mathcal{S}$ keeps $E_{\mathbf{e}}|\zeta\rangle$ the same.
Consider the completely mixed quantum state on $Q(\sigma_{J}(C))$
$\rho=\frac{1}{\dim Q(\sigma_{J}(C))}[|\zeta\rangle\langle\zeta|+\sum_{\mu\in U}|\mu\rangle\langle\mu|],$
where $\{|\zeta\rangle\cup\{|\mu\rangle:\mu\in U\}$ forms an orthonormal basis of $Q(\sigma_{J}(C))$. Since $|\zeta\rangle$ is uncorrectable, the fidelity between $\rho$ and $[\mathcal{S}(\rho\mbox{ after erasure }\mathbf{e})]$ is not one (see also [42, Exercise 9.23]). This fact, that of $\rho$ is the partial trace of $|\Phi\rangle\langle\Phi|$ over $\overline{J}$ and $\mathcal{E}$ and Equalitities (58) in [47] show that the fidelity between $|\Phi\rangle\langle\Phi|$ and $\left[\mathcal{S}\otimes\mathcal{I}_{\overline{J},\mathcal{E}}\left(|\Phi\rangle\langle\Phi|\mbox{ after erasure }\mathbf{e}\right)\right]$ is strictly less than one.
By [30, Theorem V.3], the minimum pure state fidelity for $Q(\sigma_{J}(C))$ and $\mathcal{S}$ is not one. Therefore, for every recovery procedure $\mathcal{R}^{J}_{Q(\sigma_{J}(C)),I}$, the minimum pure state fidelity neither is one. This means that there is no local recovery procedure for $J$ with erasures at $I$, which is a contradiction and it finishes the proof. ∎
Our next result proves that the requirement indicated in Lemma 13 holds whenever $C$ is symplectic self-dual and $\mathcal{E}=\mathbb{C}$. We note that given a code $C\subseteq C^{\perp_{s}}\subseteq\mathbb{F}_{p}^{2n}$, by [4, Section 20], one can always find a $p$-ary code $C_{\mathrm{max}}$ such that
(11) $C\subseteq C_{\mathrm{max}}=C_{\mathrm{max}}^{\perp_{s}}\subseteq C^{\perp_{s}}.$
Lemma 14.
Let $C$ be a symplectic self-dual $p$-ary classical code, that is $C=C^{\perp_{s}}\subseteq\mathbb{F}_{p}^{2n}$. Let $J\subsetneq\{1,\ldots,n\}$ be a nonempty set of indices and set $\overline{J}=\{1,\ldots,n\}\setminus J$. Then, for each $|\varphi\rangle\in Q(C)$, $\mathrm{tr}_{\overline{J}}[|\varphi\rangle\langle\varphi|]$ is a completely mixed state on $Q(\sigma_{J}(C))$.
Proof.
We start by constructing two symplectic self-dual codes included in $\mathbb{F}_{p}^{2\,\mathrm{card}\,(J)}$, $\mathbb{F}_{p}^{2\,(n-\mathrm{card}\,(J))}$, respectively, which we denote by $G$ and $H$. These codes will allow us to construct an auxiliary vector $|\varphi^{\prime}\rangle\in\mathbb{C}^{p^{n}}$ such that $\mathrm{tr}_{\overline{J}}[|\varphi^{\prime}\rangle\langle\varphi^{\prime}|]$ is a completely mixed state on $Q(\sigma_{J}(C))$. In the last part of the proof, we will prove our statement with the help of $|\varphi^{\prime}\rangle$.
Consider the code $\sigma_{J}(C)$ and a code $(\sigma_{J}(C))_{\mathrm{max}}$ as described in (11). Then, $\sigma_{J}(C)\subseteq(\sigma_{J}(C))_{\mathrm{max}}=((\sigma_{J}(C))_{\mathrm{max}})^{\perp_{s}}\subseteq\pi_{J}(C^{\perp_{s}})=\pi_{J}(C)$. For simplicity, denote $G:=(\sigma_{J}(C))_{\mathrm{max}}$. Define
$D_{1}:=\{\mathbf{x}\in C\;:\;\pi_{J}(\mathbf{x})\in G\}$
and, then,
(12) $\pi_{\overline{J}}(D_{1})\subseteq(\pi_{\overline{J}}(D_{1}))^{\perp_{s}}.$
Again, to simplify our notation, set $H:=\pi_{\overline{J}}(D_{1})$. To prove (12), let $\pi_{\overline{J}}(\mathbf{x}),\pi_{\overline{J}}(\mathbf{y})\in H$, then the facts that $\mathbf{x},\mathbf{y}\in C=C^{\perp_{s}}$ and $\pi_{J}(\mathbf{x}),\pi_{J}(\mathbf{y})\in G=G^{\perp_{s}}$ prove the symplectic orhogonality of $\pi_{\overline{J}}(\mathbf{x})$ and $\pi_{\overline{J}}(\mathbf{y})$, and therefore Inclusion (12).
The following sequence of equalities relates $\dim D_{1}$ to the cardinality of $\overline{J}$ and the dimension of $\sigma_{J}(C)$.
$\dim D_{1}$
$=$ $\dim C-[(\underbrace{\dim C-\dim\sigma_{\overline{J}}(C)}_{=\;\dim\pi_{J}(C)})-(\underbrace{\dim D_{1}-\dim\sigma_{\overline{J}}(C)}_{=\;\dim G}]$
$=$ $\dim C-(\underbrace{\dim\pi_{J}(C)-\dim G}_{=\;\mathrm{card}\,(J)-\dim(\sigma_{J}(C))})$
$=$ $\mathrm{card}\,(\overline{J})+\dim\sigma_{J}(C).$
Let us explain the above equalities. Set $\mathbf{0}_{J}$ the zero vector with coordinates corresponding to $J$. In the first equality, the first (respectively, second) equality under the curly brackets comes from the exact sequence $0\rightarrow\mathbf{0}_{J}\bigoplus\sigma_{\overline{J}}(C)\rightarrow C\xrightarrow{\pi_{J}}\pi_{J}(C)\rightarrow 0$ (respectively, $0\rightarrow\mathbf{0}_{J}\bigoplus\sigma_{\overline{J}}(C)\rightarrow D_{1}\xrightarrow{\pi_{J}}G\rightarrow 0$); the fact that the kernel in the second exact sequence coincides with $\mathbf{0}_{J}\bigoplus\sigma_{\overline{J}}(C)$ holds by the definition of $D_{1}$.
The equality under the curly brackets in the second equality follows, on the one hand, from the chain of equalities $[\pi_{J}(C)]^{\perp_{s}}=\sigma_{J}(C^{\perp_{s}})=\sigma_{J}(C)$ (see [43]) and from the fact that $C$ is self-dual, which proves that $\dim\pi_{J}(C)=2\,\mathrm{card}\,(J)-\dim(\sigma_{J}(C))$; and, on the other hand, from the equality $\dim(G)=\mathrm{card}\,(J)$, which is true because $G$ is self-dual.
Finally, the last equality is clear by noting that $C$ is self-dual and therefore $\dim C=\mathrm{card}\,(J)+\mathrm{card}\,(\overline{J})$.
Now, since $C$ is self-dual, one has the exact sequence $0\rightarrow\mathbf{0}_{\overline{J}}\bigoplus\sigma_{J}(C)\rightarrow D_{1}\xrightarrow{\pi_{\overline{J}}}H\rightarrow 0$ and therefore, $\dim H=\dim D_{1}-\dim\sigma_{J}(C)=\mathrm{card}\,(\overline{J})$ which, together with (12), proves that $H:=\pi_{\overline{J}}(D_{1})\subseteq\mathbb{F}_{p}^{2\,\mathrm{card}\,(\overline{J})}$ is self-dual. Recall that $G=(\sigma_{J}(C))_{\mathrm{max}}$.
Take $|\varphi_{J}\rangle\in Q(G)$ and $|\varphi_{\overline{J}}\rangle\in Q(H)$. Our past reasoning shows that $D_{1}\subseteq G\times H$ and then $|\varphi_{J}\rangle\otimes|\varphi_{\overline{J}}\rangle$ is an eigenvalue of $E_{\mathbf{y}_{1}}$ for any $\mathbf{y}_{1}\in D_{1}$. Let $D_{2}$ be an $\mathbb{F}_{p}$-linear space such that $C=D_{1}\bigoplus D_{2}$, then $\dim D_{2}=\mathrm{card}\,(J)-\dim\sigma_{J}(C)$ by the previous sequence of equalities. Now consider the column vector
(13) $|\varphi^{\prime}\rangle=\frac{1}{\sqrt{\mathrm{card}\,(D_{2})}}\sum_{\mathbf{x}\in D_{2}}(E_{\pi_{J}(\mathbf{x})}|\varphi_{J}\rangle)\otimes(E_{\pi_{\overline{J}}(\mathbf{x})}|\varphi_{\overline{J}}\rangle),$
where $E$ are operators in the corresponding error group as defined in Subsection 2.2. Any $\mathbf{y}\in C$ is expressed $\mathbf{y}=\mathbf{y}_{1}+\mathbf{y}_{2}$, with $\mathbf{y}_{i}\in D_{i}$, $i=1,2$ and then $E_{\mathbf{y}}|\varphi^{\prime}\rangle=E_{\mathbf{y}_{2}}E_{\mathbf{y}_{1}}|\varphi^{\prime}\rangle$. Now,
$E_{\mathbf{y}_{2}}|\varphi^{\prime}\rangle$
$=$ $\frac{1}{\sqrt{\mathrm{card}\,(D_{2})}}\sum_{\mathbf{x}\in D_{2}}(E_{\pi_{J}(\mathbf{x}+\mathbf{y}_{2})}|\varphi_{J}\rangle)\otimes(E_{\pi_{\overline{J}}(\mathbf{x}+\mathbf{y}_{2})}|\varphi_{\overline{J}}\rangle)$
$=$ $\frac{1}{\sqrt{\mathrm{card}\,(D_{2})}}\sum_{\mathbf{x}\in\mathbf{y}_{2}+D_{2}}(E_{\pi_{J}(\mathbf{x})}|\varphi_{J}\rangle)\otimes(E_{\pi_{\overline{J}}(\mathbf{x})}|\varphi_{\overline{J}}\rangle)$
$=$ $\frac{1}{\sqrt{\mathrm{card}\,(D_{2})}}\sum_{\mathbf{x}\in D_{2}}(E_{\pi_{J}(\mathbf{x})}|\varphi_{J}\rangle)\otimes(E_{\pi_{\overline{J}}(\mathbf{x})}|\varphi_{\overline{J}}\rangle)$
$=$ $|\varphi^{\prime}\rangle.$
Recall that $|\varphi^{\prime}\rangle$ is an eigenvector of $E_{\mathbf{y}_{1}}$ (for some eigenvalue $\lambda(E_{\mathbf{y}_{1}})$). Set $\overline{Q}(C)$ the stabilizer code having the exactly same eigenvalues as $|\varphi^{\prime}\rangle$. That is, $\overline{Q}(C)$ is a stabilizer code defined by $C$ but associated to a character determined by $|\varphi^{\prime}\rangle$. It is an $[[n,0]]_{p}$ quantum code with orthonormal basis $\{|\varphi^{\prime}\rangle\}$. Note that $Q(G)\subseteq Q(\sigma_{J}(C))$ and $Q(H)\subseteq Q(\sigma_{\overline{J}}(C))$, as $G\supseteq\sigma_{J}(C)$ and $H\supseteq\sigma_{\overline{J}}(C)$. The stabilizer codes $Q(\sigma_{J}(C))$ and $\overline{Q}(C)$ (respectively, $Q(\sigma_{\overline{J}}(C))$ and $\overline{Q}(C)$) have the same eigenvalue $\lambda(E_{\mathbf{z}})$ corresponding to $\mathbf{z}\in\sigma_{J}(C)$ (respectively, $\mathbf{z}\in\sigma_{\overline{J}}(C)$).
From our definitions, one has that $D_{1}=C\cap\pi_{J}^{-1}(G)$ and $D_{1}=C\cap\pi_{\overline{J}}^{-1}(H)$. Then, picking $\mathbf{x}\neq\mathbf{y}\in D_{2}$, one gets that $\pi_{J}(\mathbf{x})+G\neq\pi_{J}(\mathbf{y})+G$ and, analogously $\pi_{\overline{J}}(\mathbf{x})+H\neq\pi_{\overline{J}}(\mathbf{y})+H$. Therefore, $E_{\pi_{J}(\mathbf{x})}|\varphi_{J}\rangle$ (respectively, $E_{\pi_{\overline{J}}(\mathbf{x})}|\varphi_{\overline{J}}\rangle$) is orthogonal to $E_{\pi_{J}(\mathbf{y})}|\varphi_{J}\rangle$ (respectively, $E_{\pi_{\overline{J}}(\mathbf{y})}|\varphi_{\overline{J}}\rangle$). Finally, the fact that $\dim D_{2}=\mathrm{card}\,(J)-\dim\sigma_{J}(C)=\dim Q(\sigma_{J}(C))$ allows us to conclude that $\{E_{\pi_{J}(\mathbf{x})}|\varphi_{J}\rangle\;:\;\mathbf{x}\in D_{2}\}$ is an orthonormal basis of the complex linear space $Q(\sigma_{J}(C))$. Also, $\{E_{\pi_{\overline{J}}(\mathbf{x})}|\varphi_{\overline{J}}\rangle\;:\;\mathbf{x}\in D_{2}\}$ is an orthonormal subset of $Q(\sigma_{\overline{J}}(C))$.
Then, we conclude that
(14) $\mathrm{tr}_{\overline{J}}\left[|\varphi^{\prime}\rangle\langle\varphi^{\prime}|\right]=\frac{1}{\mathrm{card}\,(D_{2})}\sum_{\mathbf{x}\in D_{2}}E_{\pi_{J}(\mathbf{x})}|\varphi_{J}\rangle\langle\varphi_{J}|E_{\pi_{J}(\mathbf{x})}^{\dagger},$
which is a completely mixed state on $Q(\sigma_{J}(C))$.
Finally, consider $|\varphi\rangle$ as in the statement and we are going to prove that $\mathrm{tr}_{\overline{J}}\left[|\varphi\rangle\langle\varphi|\right]$ is a completely mixed state on $Q(\sigma_{J}(C))$. The complex linear space $Q(\sigma_{J}(C))\otimes Q(\sigma_{\overline{J}}(C))$ corresponds to the code $\sigma_{J}(C)\times\sigma_{\overline{J}}(C)$ whose symplectic dual is $\pi_{J}(C)\times\pi_{\overline{J}}(C)$. Bearing in mind that, in a stabilizer code, elements $E_{\mathbf{y}}$ modify eigenspaces if and only if $\mathbf{y}$ is not in the symplectic dual of the defining classical code, we deduce that the equality of linear spaces $E_{\mathbf{y}}[Q(\sigma_{J}(C))\otimes Q(\sigma_{\overline{J}}(C))]=Q(\sigma_{J}(C))\otimes Q(\sigma_{\overline{J}}(C))$ is true if, and only if, $\mathbf{y}\in\pi_{J}(C)\times\pi_{\overline{J}}(C)$. Now, by Equality (14), the above introduced stabilizer code $\overline{Q}(C)$ satisfies $\overline{Q}(C)\subseteq Q(\sigma_{J}(C))\otimes Q(\sigma_{\overline{J}}(C))$ and by Proposition 11, $Q(C)\subseteq Q(\sigma_{J}(C))\otimes Q(\sigma_{\overline{J}}(C))$. Thus, there exists $\mathbf{x}\in\pi_{J}(C)\times\pi_{\overline{J}}(C)\setminus C$ satisfying the three following equalities: $E_{\mathbf{x}}(\overline{Q}(C))=Q(C)$, $E_{\pi_{J}(\mathbf{x})}(Q(\sigma_{J}(C))=Q(\sigma_{J}(C))$ and $E_{\pi_{\overline{J}}(\mathbf{x})}(Q(\sigma_{\overline{J}}(C))=Q(\sigma_{\overline{J}}(C))$. Therefore, $E_{\mathbf{x}}|\varphi^{\prime}\rangle$ is a scalar multiple of the vector $|\varphi\rangle\in Q(C)$. This concludes the proof because the fact that the expression in (14) is a completely mixed state on $Q(\sigma_{J}(C))$ implies that the same happens for $\mathrm{tr}_{\overline{J}}\left[|\varphi\rangle\langle\varphi|\right]$. ∎
Now we state and prove the last of the aforementioned lemmas. Remind that we desire to show that Equality (10) is a necessary condition for $Q(C)$ to be an $(I,J)$-locally recoverable code.
Lemma 15.
Let $Q(C)\subseteq\mathbb{C}^{p^{n}}$ be a stabilizer code given by a symplectic self-orthogonal $p$-ary code $C\subseteq\mathbb{F}_{p}^{2n}$, and consider sets of indices $\emptyset\neq I\subsetneq J\varsubsetneq\{1,\ldots,n\}$. Set $\overline{J}=\{1,\ldots,n\}\setminus J$. Then, there exists an environment $\mathcal{E}$ surrounding $\mathbb{C}^{p^{n}}$ and an entangled quantum codeword $|\Phi\rangle\in Q(C)\otimes\mathcal{E}$ such that the partial trace of $|\Phi\rangle\langle\Phi|$ over $\overline{J}$ and $\mathcal{E}$ is a completely mixed state on $Q(\sigma(C))$.
Proof.
Consider the code $C_{\mathrm{max}}$ defined in (11) and its shortening $\sigma_{J}(C_{\mathrm{max}})$. Then
$Q(\sigma_{J}(C_{\mathrm{max}}))\subseteq Q(\sigma_{J}(C)).$
Furthermore, $Q(\sigma_{J}(C_{\mathrm{max}}))$ is orthogonal to $E_{\mathbf{z}}Q(\sigma_{J}(C_{\mathrm{max}}))$ if and only if $\mathbf{z}\not\in[\sigma_{J}(C_{\mathrm{max}})]^{\perp_{s}}$.
When $\mathbf{z}$ runs over $\pi_{J}(C^{\perp_{s}})=[\sigma_{J}(C)]^{\perp_{s}}$, the cardinality of the quotient linear space $[(\pi_{J}(C^{\perp_{s}})/(\sigma_{J}(C_{\mathrm{max}}))^{\perp_{s}}]$ coincides with the number of different spaces $E_{\mathbf{z}}Q(\sigma_{J}(C_{\mathrm{max}}))$. Then, one gets the following chain of equalities:
$\dim\sum_{\mathbf{z}\in\pi_{J}(C^{\perp_{s}})}E_{\mathbf{z}}Q(\sigma_{J}(C_{\mathrm{max}}))$ $=$ $\mathrm{card}\,[(\pi_{J}(C^{\perp_{s}})/(\sigma_{J}(C_{\mathrm{max}}))^{\perp_{s}}]\dim Q(\sigma_{J}(C_{\mathrm{max}}))$
$=$ $p^{\dim(\sigma_{J}(C))^{\perp_{s}}-\dim(\sigma_{J}(C_{\mathrm{max}}))^{\perp_{s}}}\times p^{\mathrm{card}\,J-\dim\sigma_{J}(C_{\mathrm{max}})}$
$=$ $p^{\dim(\sigma_{J}(C_{\mathrm{max}}))-\dim\sigma_{J}(C)}\times p^{\mathrm{card}\,J-\dim\sigma_{J}(C_{\mathrm{max}})}$
$=$ $p^{\mathrm{card}\,J-\dim\sigma_{J}(C)}=\dim Q(\sigma_{J}(C)).$
The equality of dimensions we have just obtained and the inclusion
$\sum_{\mathbf{z}\in\pi_{J}(C^{\perp_{s}})}E_{\mathbf{z}}Q(\sigma_{J}(C_{\mathrm{max}}))\subseteq Q(\sigma_{J}(C)),$
prove that
(15) $\sum_{\mathbf{z}\in\pi_{J}(C^{\perp_{s}})}E_{\mathbf{z}}Q(\sigma_{J}(C_{\mathrm{max}}))=Q(\sigma_{J}(C)).$
Recall that $(\sigma_{J}(C_{\mathrm{max}}))^{\perp_{s}}=\pi_{J}(C_{\mathrm{max}})$. Consider a set $T_{J}\subseteq\mathbf{F}_{p}^{2\mathrm{card}\,(J)}$ of representatives of the cosets in the quotient $[\pi_{J}(C^{\perp_{s}})/\pi_{J}(C_{\mathrm{max}})]$. Since $\mathbf{z}\in(\sigma_{J}(C_{\mathrm{max}}))^{\perp_{s}}$ implies the equality of spaces $Q(\sigma_{J}(C_{\mathrm{max}}))=E_{\mathbf{z}}Q(\sigma_{J}(C_{\mathrm{max}}))$, Equation (15) can be rewritten as
(16) $\bigoplus_{\mathbf{z}\in T_{J}}E_{\mathbf{z}}Q(\sigma_{J}(C_{\mathrm{max}}))=Q(\sigma_{J}(C)).$
Now, on the one hand, choose $T\subseteq C^{\perp_{s}}$ such that $\pi_{J}(T)=T_{J}$ and $\mathrm{card}\,(T)=\mathrm{card}\,(T_{J})=p^{\dim\pi_{J}(C^{\perp_{s}})-\dim\pi_{J}(C_{\mathrm{max}})}$. And, on the other hand, consider an environment $\mathcal{E}$ (surrounding $\mathbb{C}^{p^{n}}$) with an orthonormal basis $\{|\mathbf{y}\rangle_{\mathcal{E}}\;:\;\mathbf{y}\in T\}$. $T$, $\mathcal{E}$ and any fixed $|\varphi\rangle\in Q(C_{\mathrm{max}})\subseteq Q(C)$ allow us to construct the announced vector $|\Phi\rangle$:
(17) $|\Phi\rangle=\frac{1}{\sqrt{\mathrm{card}\,(T)}}\sum_{\mathbf{y}\in T}E_{\mathbf{y}}|\varphi\rangle\otimes|\mathbf{y}\rangle_{\mathcal{E}}.$
Equality (16) and Lemma 14 prove that the partial trace of $|\Phi\rangle\langle\Phi|$ over $\overline{J}$ and $\mathcal{E}$ is a completely mixed state on $Q(\sigma_{J}(C))$. This concludes the proof. ∎
Lemmas 15 and 13 prove our before announced result on the necessity of Condition (10) for $(I,J)$-local recoverability. Let us state it.
Proposition 16.
Let $Q(C)\subseteq\mathbb{C}^{p^{n}}$ be a stabilizer code given by a symplectic self-orthogonal $p$-ary code $C\subseteq\mathbb{F}_{p}^{2n}$ and consider sets of indices $\emptyset\neq I\subsetneq J\subsetneq\{1,\ldots,n\}$. The codes’ equality stated in (10):
$\sigma_{I}\left[\pi_{J}\left(C^{\perp_{s}}\right)\right]=\sigma_{I}(C)$
is a necessary condition for $Q(C)$ to be $(I,J)$-locally recoverable.
We conclude with our main result in this subsection. It follows from Theorem 10 and, by reasoning as in [21, Section 2.1], from Propositions 12 and 16.
Theorem 17.
Let $Q(C)\subseteq\mathbb{C}^{q^{n}}$ be a stabilizer code given by a symplectic self-orthogonal $q$-ary code $C\subseteq\mathbb{F}_{q}^{2n}$, $q$ being a prime power. Consider sets of indices $I$ and $J$ such that $\emptyset\neq I\subsetneq J\subseteq\{1,\ldots,n\}$. Then, $Q(C)$ is $(I,J)$-locally recoverable if and only if Condition (10):
$\sigma_{I}\left[\pi_{J}\left(C^{\perp_{s}}\right)\right]=\sigma_{I}(C)$
holds.
3.2.3. Generalized symplectic weight and $(I,J)$-local recoverability
This subsection recalls the concept of generalized symplectic weight and relates it with (quantum) $(I,J)$-local recoverability. Let us start with the definition of generalized symplectic weight [40, Definition 15].
Definition 18.
Let $C\subseteq\mathbb{F}_{q}^{2n}$ an $\mathbb{F}_{q}$-linear space. For each index $1\leq t\leq\dim C$, the $t$th generalized symplectic weight of $C$, $\mathrm{gsw}_{t}(C)$ is defined as
$\mathrm{gsw}_{t}(C):=\min\{\mathrm{card}\,(J):\dim(\sigma_{J}(C))\geq t\}.$
Note that $\mathrm{gsw}_{1}(C)=\mathrm{swt}(C)$, $\mathrm{gsw}_{t}(C)$ needs not be equal to $\min\{\mathrm{card}\,(J):\dim(\sigma_{J}(C)=t\}$ and $\mathrm{gsw}_{t+1}(C)\geq\mathrm{gsw}_{t}(C)$, $11\leq t\leq\dim C-$ 1.
Our first result is the following one.
Proposition 19.
Let $Q(C)$ be a stabilizer code defined by a symplectic self-orthogonal code $C\subseteq\mathbb{F}_{q}^{2n}$. Consider sets of indices $\emptyset\neq I\subsetneq J\subseteq\{1,\ldots,n\}$ and assume that $\mathrm{card}\,(J)\geq n-\mathrm{swt}(C^{\perp_{s}})+\mathrm{card}\,(I)+1$, then $Q(C)$ is $(I,J)$-locally recoverable.
Proof.
The fact that $\mathrm{swt}(C^{\perp_{s}})\geq n-\mathrm{card}\,(J)+\mathrm{card}\,(I)+1$ implies that
$\mathrm{swt}[\pi_{J}(C^{\perp_{s}})]\geq\mathrm{card}\,(I)+1,$
and then $\sigma_{I}[\pi_{J}(C^{\perp_{s}})]=\{\mathbf{0}\}$. Thus, Equality (10) holds and the result follows by Theorem 17. ∎
Next, we give our last result in this subsection. It provides two inequalities that must hold when $Q(C)$ is $(I,J)$-locally recoverable. We state the counter-reciprocal statement.
Proposition 20.
Keep the notation as in Proposition 19, $Q(C)$ being an $[[n,k]]_{q}$ stabilizer code. Assume that $\mathrm{swt}(C)\geq\mathrm{card}\,(I)+1$. Then, Equality (10) is false whenever the following two inequalities hold:
(18) $n+k-2\,\mathrm{card}\,(J)+2\,\mathrm{card}\,(I)>0$
and
(19) $\mathrm{gsw}_{n+k-2\,\mathrm{card}\,(J)+2\,\mathrm{card}\,(I)}(C^{\perp_{s}})\geq n-\mathrm{card}\,(J)+1.$
Proof.
The assumption $\mathrm{swt}(C)\geq\mathrm{card}\,(I)+1$ implies that the right hand side of Equality (10) vanishes.
The shortening at $I$ of $\pi_{J}\left(C^{\perp_{s}}\right)$ removes $(2\,\mathrm{card}\,(J)-2\,\mathrm{card}\,(I))$ components and then
(20) $\dim\pi_{J}\left(C^{\perp_{s}}\right)\geq 2\,\mathrm{card}\,(J)-2\,\mathrm{card}\,(I)+1$
implies that $\sigma_{I}\left[\pi_{J}\left(C^{\perp_{s}}\right)\right]\neq\{\mathbf{0}\}$ and therefore Equality (10) is not true.
Let us show that Inequalities (18) and (19) imply (20).
Consider the linear map $\pi_{J}:\mathbb{F}_{q}^{2n}\rightarrow\mathbb{F}_{q}^{2\,\mathrm{card}\,(J)}$, whose kernel is $\sigma_{\overline{J}}(\mathbb{F}_{q}^{2n})$ and thus
$\dim\pi_{J}(C^{\perp_{s}})=n+k-\dim(C^{\perp_{s}}\cap\sigma_{\overline{J}}(\mathbf{F}_{q}^{2n}))=n+k-\dim(\sigma_{\overline{J}}(C^{\perp_{s}})).$
Then, Inequality (20) is equivalent to the inequality
$n+k-\dim(\sigma_{\overline{J}}(C^{\perp_{s}}))>2\,\mathrm{card}\,(J)-2\,\mathrm{card}\,(I)$
and this inequality is equivalent to
(21) $\dim(\sigma_{\overline{J}}(C^{\perp_{s}}))<n+k-2\,\mathrm{card}\,(J)+2\,\mathrm{card}\,(I).$
The fact that inequalities (18) and (19) imply (21) concludes the proof. ∎
Example 1.
Contrary to what happens with classical error-correcting codes, the above inequality $\mathrm{gsw}_{t+1}(C)\geq\mathrm{gsw}_{t}(C)$ needs not be strict. Indeed, Let $C_{H}^{\perp_{e}}$ be the $[7,4,3]_{2}$ Hamming code with $C_{H}\subseteq C_{H}^{\perp_{e}}$. By [55], the generalized Hamming weight hierarchy of $C_{H}$ is $(4,6,7)$ and that of $C_{H}^{\perp_{e}}$ is $(3,5,6,7)$. Consider the code $C:=C_{H}\times C_{H}\subseteq\mathbb{F}_{2}^{14}$, which defines the $[[7,1,3]]_{2}$ Steane quantum stabilizer code. We have $\mathrm{gsw}_{1}(C^{\perp_{s}})=\mathrm{gsw}_{2}(C^{\perp_{s}})=3$ and $\mathrm{gsw}_{3}(C^{\perp_{s}})=\mathrm{gsw}_{4}(C^{\perp_{s}})=5$.
In addition, Proposition 19 proves that for any $i\in J$ with $\mathrm{card}\,(J)=6$, an erasure at $i$ can be recovered by $J$.
3.3. Quantum $(r,\delta)$-local recoverability
This brief subsection takes the previous results to the language of quantum $(r,\delta)$-locally recoverable codes. Our results follow directly from preceding results in this section.
The first one follows straightforwardly from Theorem 17.
Theorem 21.
Let $Q(C)\subseteq\mathbb{C}^{q^{n}}$ be a stabilizer code given by a symplectic self-orthogonal $q$-ary linear code $C\subseteq\mathbb{F}_{q}^{2n}$, $q$ being a prime power. $Q(C)$ is a quantum $(r,\delta)$-LRC if and only if for all index $i\in\{1,\ldots,n\}$, there exists a set $J\ni i$ of cardinality $\leq r+\delta-1$ such that the following equality holds
$\sigma_{I}\left[\pi_{J}\left(C^{\perp_{s}}\right)\right]=\sigma_{I}(C)$
for any subset $I\subsetneq J$ of cardinality $\delta-1$.
Our second result can be deduced from Propositions 19 and 20.
Proposition 22.
Let $Q(C)$ be an stabilizer code defined by a symplectic self-orthogonal linear code $C\subseteq\mathbb{F}_{q}^{2n}$.
  1. (1)
    Suppose that $\mathrm{swt}(C^{\perp_{s}})\geq n-r+1$. Then, for all possible values $\delta$, $Q(C)$ is a quantum $(r,\delta)$-LRC.
  2. (2)
    Assume that $\mathrm{swt}(C)\geq\delta$. If $Q(C)$ is an $[[n,k]]_{q}$ quantum $(r,\delta)$-LRC, then either $n+k\leq 2r$ or $n+k>2r$ and $\mathrm{gsw}_{n+k-2r}(C^{\perp_{s}})<n-(r+\delta)+2$.
4. Quantum $(r,\delta)$-LRCs from Hermitian and Euclidean inner products. Relation to classical $(r,\delta)$-LRCs. A Singleton-like bound
We divide this section in three parts. The first one gives conditions for local recoverability of quantum codes defined by Hermitian or Euclidean self-orthogonal codes. In the second one, we relate classical and quantum local recoverability while a Singleton-like bound for quantum $(r,\delta)$-LRCs coming from Hermitian and Euclidean dual-containing codes is stated in the last part.
4.1. Quantum $(r,\delta)$-LRCs from Hermitian and Euclidean inner products
As we said before, quantum stabilizer codes can be constructed by using Hermitian or Euclidean inner products instead of symplectic forms. In these cases, Theorems 17 and 21 can be straightforwardly translated.
Recall that the Hermitian product of two vectors $\mathbf{x},\mathbf{y}$ in $\mathbb{F}_{q^{2}}^{n}$ is defined as $\mathbf{x}\cdot_{h}\mathbf{y}=\mathbf{x}^{q}\cdot_{e}\mathbf{y}$, where $\mathbf{x}^{q}$ is the componentwise $q$th-power of $\mathbf{x}$ and $\cdot_{e}$ stands for the Euclidean inner product. Also, take a normal basis $\{\varsigma,\varsigma^{q}\}$ of $\mathbb{F}_{q^{2}}$ over $\mathbb{F}_{q}$ and remember that the trace-alternating product of the previous vectors $\mathbf{x},\mathbf{y}$ is defined as $\mathbf{x}\cdot_{a}\mathbf{y}:=\mathrm{tr}_{q/p}\left(\frac{\mathbf{x}\cdot_{e}\mathbf{y}^{q}-\mathbf{x}^{q}\cdot_{e}\mathbf{y}}{\varsigma^{2q}-\varsigma^{2}}\right)$. Consider the one to one isometry $\psi:\mathbb{F}_{q}^{2n}\rightarrow\mathbb{F}_{q^{2}}^{n}$ defined by $\psi(\mathbf{a}|\mathbf{b})=\varsigma\mathbf{a}+\varsigma^{q}\mathbf{b}$ and then
$(\mathbf{a}|\mathbf{b})._{\mathrm{trs}}(\mathbf{c}|\mathbf{d})=\psi\left((\mathbf{a}|\mathbf{b})\right)\cdot_{a}\psi\left((\mathbf{c}|\mathbf{d})\right).$
By [29, Lemma 18], an $\mathbb{F}_{q^{2}}$-linear subspace $D$ of $\mathbb{F}_{q^{2}}^{n}$ satisfies $D^{\perp_{a}}=D^{\perp_{h}}$ and, by [29, Theorem 15], the existence of a quantum stabilizer code is equivalent to that of an additive code $D$ in $\mathbb{F}_{q^{2}}^{n}$ satisfying $D\subseteq D^{\perp_{a}}$. These results allow us to translate ours on local recoverability to stabilizer codes defined by self-orthogonal, with respect to the Hermitian inner product, linear subspaces of $\mathbb{F}_{q^{2}}^{n}$.
The same happens when considering $\mathbb{F}_{q}$-linear subspaces $C_{1}$ and $C_{2}$ included in $\mathbb{F}_{q}^{n}$ and such that $C_{2}\subseteq C_{1}^{\perp_{e}}$. Note that, in this case, one should consider $C_{1}\times C_{2}$ and, by the CSS procedure, use the symplectic construction.
Let us state our results.
Theorem 23.
Let $Q(C)\subseteq\mathbb{C}^{q^{n}}$ be a stabilizer code given by a $q^{2}$-ary (respectively, $q$-ary) Hermitian (respectively, Euclidean) self-orthogonal code $C\subseteq\mathbb{F}_{q^{2}}^{n}$ (respectively, $C\subseteq\mathbb{F}_{q}^{n}$), $q$ being a prime power. Consider sets of indices $I$ and $J$ such that $\emptyset\neq I\subsetneq J\subseteq\{1,\ldots,n\}$. Then, $Q(C)$ is $(I,J)$-locally recoverable if and only if the equality
(22) $\sigma_{I}\left[\pi_{J}\left(C^{\perp_{h}}\right)\right]=\sigma_{I}(C)$
(respectively,
(23) $\sigma_{I}\left[\pi_{J}\left(C^{\perp_{e}}\right)\right]=\sigma_{I}(C)\,)$
holds.
Theorem 24.
Let $Q(C)\subseteq\mathbb{C}^{q^{n}}$ be a stabilizer code given by a $q^{2}$-ary (respectively $q$-ary) Hermitian (respectively, Euclidean) self-orthogonal code $C\subseteq\mathbb{F}_{q^{2}}^{n}$ (respectively, $C\subseteq\mathbb{F}_{q}^{n}$), $q$ being a prime power. $Q(C)$ is a quantum $(r,\delta)$-LRC if and only if, for all index $i\in\{1,\ldots,n\}$, there exists a set $J\ni i$ of cardinality $\leq r+\delta-1$ such that Equality (22) (respectively, Equality (23)) holds for any subset $I\subsetneq J$ of cardinality $\delta-1$.
Theorem 23 contemplates a unique code $C$ in the CSS case. For the general CSS construction we state the following result.
Proposition 25.
Let $Q(C)\subseteq\mathbb{C}^{q^{n}}$ be a stabilizer code given by two codes $C_{1}$ and $C_{2}$ included in $\mathbb{F}_{q}^{n}$ and such that $C_{2}\subseteq C_{1}^{\perp_{e}}$. Then, $Q(C)$ is $(I,J)$-locally recoverable if and only if the following equalities
$\sigma_{I}\left[\pi_{J}\left(C_{1}^{\perp_{e}}\right)\right]=\sigma_{I}(C_{2})$
and
$\sigma_{I}\left[\pi_{J}\left(C_{2}^{\perp_{e}}\right)\right]=\sigma_{I}(C_{1})$
hold.
Proof.
By Theorem 17, $Q(C)$ is $(I,J)$-locally recoverable if and only if
$\sigma_{I}\left[\pi_{J}\left(\left(C_{1}\times C_{2}\right)^{\perp_{s}}\right)\right]=\sigma_{I}[(C_{1}\times C_{2})]$
which is equivalent to
$\sigma_{I}\left[\pi_{J}\left(C_{2}^{\perp_{e}}\right)\right]\times\sigma_{I}\left[\pi_{J}\left(C_{1}^{\perp_{e}}\right)\right]=\sigma_{I}(C_{1})\times\sigma_{I}(C_{2}),$
which is equivalent to the conditions in the statement. This ends the proof. ∎
4.2. Quantum and classical $(r,\delta)$-LRCs
We start by stating and proving the following result concerning $(I,J)$-local recoverability of classical codes.
Proposition 26.
Let $C$ be a $q^{2}$-ary linear code included in $\mathbb{F}_{q^{2}}^{n}$ (respectively, a $q$-ary linear code included in $\mathbb{F}_{q}^{n}$). Consider sets of indices $I$ and $J$ such that $\emptyset\neq I\subsetneq J\subseteq\{1,\ldots,n\}$. Then, erasures of the code $C$ at $I$ can be classically corrected from the knowledge of the coordinates at $J\setminus I$ if and only if the following equivalent conditions hold:
(24) $\sigma_{I}\left[\pi_{J}(C)\right]=\{\mathbf{0}\}\Longleftrightarrow\sigma_{I}\left[\left(\sigma_{J}(C^{\perp_{h}})\right)^{\perp_{h}}\right]=\{\mathbf{0}\}$
(respectively,
(25) $\sigma_{I}\left[\pi_{J}(C)\right]=\{\mathbf{0}\}\Longleftrightarrow\sigma_{I}\left[\left(\sigma_{J}(C^{\perp_{e}})\right)^{\perp_{e}}\right]=\{\mathbf{0}\}\,).$
Proof.
We give a proof for the Euclidean case; the Hermitian case follows analogously. Suppose that $C$ is an $[n,k]_{q}$ code. Take a basis $\{\mathbf{h}_{1},\ldots,\mathbf{h}_{n-k}\}$ of the dual space $C^{\perp_{e}}$ and consider the map $f:\mathbb{F}_{q}^{n}\rightarrow\mathbb{F}_{q}^{n-k}$ defined as $f(\mathbf{x})=(\mathbf{x}\cdot_{e}\mathbf{h}_{1},\ldots,\mathbf{x}\cdot_{e}\mathbf{h}_{n-k})$. Correction of erasures consists of solving a system of linear equations $f(\mathbf{e})=\mathbf{s}$ where $\mathbf{e}=(e_{1},\ldots,e_{n})$ is unknown and $\mathbf{s}=(s_{1},\ldots,s_{n-k})$ a syndrome obtained from the received word. Since erasures are in $I$, we have $e_{i}=0$ whenever $i\notin I$ and noticing that $\ker(f)=C$, we get a unique solution if and only if $\sigma_{I}(C)=\{\mathbf{0}\}$. Finally, we only use coordinates at $J$ for local recovery, which means that we use $\pi_{J}(C)$ instead of $C$ and the result is proved. ∎
Our next result relates $(I,J)$-local recoverability of quantum and classical codes. We assume that the classical codes satisfy dual-containing instead of self-orthogonality because, under this assumption, distances of quantum codes fit better with those of classical codes. Recall that, as we said in the introduction, we denote by $Q^{\prime}(C)$ the quantum stabilizer code given by a Hermitian or Euclidean dual-containing code $C$. Note that $Q^{\prime}(C)=Q(C^{\perp})$, where $C^{\perp}$ is the corresponding Hermitian or Euclidean dual code of $C$.
Proposition 27.
Let $C$ be a $q^{2}$-ary linear code included in $\mathbb{F}_{q^{2}}^{n}$ (respectively, a $q$-ary linear code included in $\mathbb{F}_{q}^{n}$). Assume that $C^{\perp_{h}}\subseteq C$ (respectively, $C^{\perp_{e}}\subseteq C$) and that $\dim C=\frac{n+k}{2}$. Consider sets of indices $I$ and $J$ such that $\emptyset\neq I\subsetneq J\subseteq\{1,\ldots,n\}$. Assume that $\mathrm{card}\,(I)\leq d(C^{\perp_{h}})-1$ (respectively, $\mathrm{card}\,(I)\leq d(C^{\perp_{e}})-1$).
Then, $Q^{\prime}(C)$ is an $(I,J)$-locally recoverable code if and only if erasures of the code $C$ at $I$ can be classically corrected from the knowledge of the coordinates of $C$ at $J\setminus I$.
Proof.
The assumption $\mathrm{card}\,(I)\leq d(C^{\perp_{h}})-1$ (respectively, $\mathrm{card}\,(I)\leq d(C^{\perp_{e}})-1$) implies $\sigma_{I}\left(C^{\perp_{h}}\right)=\{\mathbf{0}\}$ (respectively, $\sigma_{I}\left(C^{\perp_{e}}\right)=\{\mathbf{0}\}$. Then, noticing that we are considering dual-containing and the projection $\pi_{J}$, Equality (24) (respectively, Equality (25)) is equivalent to Equality (22) (respectively, Equality (23)), which concludes the proof. ∎
We finish this subsection with our main result relating quantum and classical $(r,\delta)$-locally recoverable codes. It follows from Proposition 27.
Theorem 28.
Let $C$ be a $q^{2}$-ary linear code included in $\mathbb{F}_{q^{2}}^{n}$ (respectively, a $q$-ary linear code included in $\mathbb{F}_{q}^{n}$). Assume that $C$ is Hermitian (respectively, Euclidean) dual-containing. Suppose also that $\delta\leq d(C^{\perp_{h}})$ (respectively, $d(C^{\perp_{e}}))$. Then, $Q^{\prime}(C)$ is a quantum $(r,\delta)$-LRC if and only if $C$ is a classical $(r,\delta)$-LRC.
Remark 29.
Corollary 3 in [37] proves that an Euclidean dual classical LRC of locality $r$ provides a quantum LRC of locality $r$.
4.3. A Singleton-like bound
Our last subsection provides a Singleton-like bound for quantum $(r,\delta)$-LRCs coming from Hermitian or Euclidean dual-containing classical codes. Our bound is proved under a suitable bound for the value $\delta$ and we deduce it from Proposition 4.
Theorem 30.
Let $C$ be a $q^{2}$-ary linear code included in $\mathbb{F}_{q^{2}}^{n}$ (respectively, a $q$-ary linear code included in $\mathbb{F}_{q}^{n}$). Assume that $C$ is Hermitian (respectively, Euclidean) dual-containing, $\dim C=\frac{n+k}{2}$ and $C$ is an $(r,\delta)$-LRC. Suppose also that $\delta\leq d(C^{\perp_{h}})$ (respectively, $\delta\leq d(C^{\perp_{e}})$. Then, the quantum $(r,\delta)$-LRC, $Q^{\prime}(C)$, has parameters $[[n,k,\geq d(C)]]_{q}$, which satisfy
(26) $k+2d(C)+2\left(\left\lceil\frac{n+k}{2r}\right\rceil-1\right)(\delta-1)\leq n+2.$
Proof.
Theorem 28 proves that $Q^{\prime}(C)$ is a quantum $(r,\delta)$-LRC. $C$ is a classical $(r,\delta)$-LRC and then, their parameters satisfy Inequality (4). Taking into account that $\dim C=\frac{n+k}{2}$, we get the inequality
$\frac{n+k}{2}+d(C)+\left(\left\lceil\frac{n+k}{2r}\right\rceil-1\right)(\delta-1)\leq n+1,$
which is equivalent to (26) and, thus, the result is proved. ∎
Definition 31.
A quantum $(r,\delta)$-LRC as before is said to be optimal whenever its parameters and $(r,\delta)$-locality attain the bound in (26).
5. Examples
We divide this section in two subsections. The first one gives some results on and examples of optimal $(r,\delta)$-LRCs, $C$, satisfying the conditions in Theorem 28 relative to Euclidean duality. Then, the stabilizer codes $Q^{\prime}(C)$ are optimal quantum $(r,\delta)$-LRCs. Our second subsection does the same but in the case of considering Hermitian duality.
5.1. Some examples of optimal quantum $(r,\delta)$-LRCs by using Euclidean duality
In this subsection, we are going to consider some families of $\emptyset$-affine variety codes to get optimal LRCs satisfying the Euclidean requirements in Theorem 28.
Recall that $q$ is a power of a prime number $p$ and consider two integer numbers $n_{1}$ and $n_{2}$ such that both $n_{1}-1$ and $n_{2}-1$ divide $q-1$. Let $\mathcal{J}$ be the ideal of the polynomial ring $\mathbb{F}_{q}[X,Y]$ generated by the binomials $X^{n_{1}}-X$ and $Y^{n_{2}}-Y$. Set
$P=\{\boldsymbol{\alpha}_{1},\dots,\boldsymbol{\alpha}_{n}\}\subseteq\mathbb{F}_{q}^{2}$
the zero-set of $\mathcal{J}$.
Let
$E=\{0,1,\dots,n_{1}-1\}\times\{0,1,\dots,n_{2}-1\}$
be the set of pairs corresponding to exponents of the monomials whose classes we are going to evaluate with the following linear evaluation map:
$\mathrm{ev}_{P}:\frac{\mathbb{F}_{q}[X,Y]}{\mathcal{J}}\rightarrow\mathbb{F}_{q}^{n}\textrm{, }\quad\mathrm{ev}_{P}(f)=\left(f(\boldsymbol{\alpha}_{1}),\dots,f(\boldsymbol{\alpha}_{n})\right).$
Then, for each set $\emptyset\neq\Delta\subseteq E$, we consider the $\emptyset$-affine variety code
$C_{\Delta}^{P}:=\langle\mathrm{ev}_{P}(X^{e_{1}}Y^{e_{2}})\;:\;(e_{1},\,e_{2})\in\Delta\rangle\subseteq\mathbb{F}_{q}^{n},$
where $\langle W\rangle$ means the $\mathbb{F}_{q}$-linear space generated by the elements in $W\subseteq\mathbb{F}_{q}^{n}$. $\emptyset$-affine variety codes were introduced in [18] and $C_{\Delta}^{P}$ is a monomial-Cartesian code as indicated in [20].
Let us show our results.
Proposition 32.
Keep the above notation and assume that $p$ divides $n_{1}$ and $n_{2}$. For each pair $(i,j)\in E$, consider the set
$\Delta=\Delta_{i,j}:=\left\{(e_{1},e_{2})\mid 0\leq e_{1}\leq i\textrm{, }0\leq e_{2}\leq j\right\}\subseteq E=\{0,\dots,n_{1}-1\}\times\{0,\dots,n_{2}-1\}.$
Then, the code $C_{\Delta}^{P}$ is an optimal $(r,\delta)$-LRC and satisfies the Euclidean conditions in Theorem 28 —giving rise to an optimal quantum $(r,\delta)$-LRC, $Q^{\prime}(C_{\Delta}^{P})$,— if, and only if, one of the following conditions hold:
  • $i\geq\frac{n_{1}-1}{2}$ and $j=n_{2}-1$, in which case $(r,\delta)=(i+1,n_{1}-i)$ (see Figure 1 (A)).
  • $i=n_{1}-1$ and $j\geq\frac{n_{2}-1}{2}$, in which case $(r,\delta)=(j+1,n_{2}-j)$ (see Figure 1 (B)).
The parameters of $Q^{\prime}(C_{\Delta}^{P})$ are $[[n_{1}n_{2},2(i+1)(j+1)-n_{1}n_{2},\geq(n_{1}-i)(n_{2}-j)]]_{q}$.
Proof.
Let us prove the first item. The second one holds by symmetry. The fact that $C_{\Delta}^{P}$ is an optimal $(r,\delta)$-LRC follows from [20, Proposition 4.1]. Now, [18, Proposition 2.2] proves that, under the conditions in the statement, $(C_{\Delta}^{P})^{\perp_{e}}\subseteq C_{\Delta}^{P}$. Finally, $i\geq\frac{n_{1}-1}{2}$ shows that
$\delta=n_{1}-i\leq i+2=d(C_{\Delta_{i,j}}^{P})^{\perp_{e}},$
where the last equality follows from [20, Remark 3.8]. Then, all the requirements in Theorem 28 are satisfied and $Q^{\prime}(C_{\Delta}^{P})$ is an optimal quantum $(r,\delta)$-LRC. The parameters of $Q^{\prime}(C_{\Delta}^{P})$ follow from Item (2) in Theorem 7. ∎
[…]
(a) Sets $\Delta_{i,n_{2}-1}$
[…]
(b) Sets $\Delta_{n_{1}-1,j}$
Figure 1. Sets $\Delta_{i,n_{2}-1}$ and $\Delta_{n_{1}-1,j}$ in Proposition 32
Our second result can be proved with the same arguments we used to prove Proposition 32, but, instead of Proposition 4.1, we need to use Proposition 4.2 in [20].
Proposition 33.
Keep the same notation and assumption as in Proposition 32. Consider the set
$\Delta=\Delta_{i,s}^{2}:=\left\{(e_{1},e_{2})\;:\;0\leq e_{1}\leq i\textrm{, }0\leq e_{2}\leq n_{2}-2\right\}\cup\left\{(e_{1},n_{2}-1)\mid 0\leq e_{1}\leq s\right\},$
where $\frac{n_{1}-2}{2}<i\leq n_{1}-2$ and $s\geq n_{1}-2-i$ (see Figure 2 (A)).
Then, the code $C_{\Delta}^{P}$ is an optimal $(r,\delta)=(i+1,n_{1}-i)$-LRC and satisfies the Euclidean conditions in Theorem 28. Therefore, it gives rise to an optimal quantum $(r,\delta)=(i+1,n_{1}-i)$-LRC, $Q^{\prime}(C_{\Delta}^{P})$.
The parameters of $Q^{\prime}(C_{\Delta}^{P})$ are $[[n_{1}n_{2},2[(i+1)(n_{2}-1)+s+1]-n_{1}n_{2},\geq(n_{1}-s)]]_{q}$.
Analogously, the code $C_{\Delta}^{P}$, where
$\Delta=\Delta_{j,s}^{2,\sigma}:=\left\{(e_{1},e_{2})\;:\;0\leq e_{1}\leq n_{1}-2\textrm{, }0\leq e_{2}\leq j\right\}\cup\left\{(n_{1}-1,e_{2})\mid 0\leq e_{2}\leq s\right\},$
for $\frac{n_{2}-2}{2}<j\leq n_{2}-2$ and $s\geq n_{2}-2-j$ (see Figure 2 (B)) is an optimal $(r,\delta)=(j+1,n_{2}-j)$-LRC, which satisfies the Euclidean conditions in Theorem 28. Therefore, it gives rise to an optimal quantum $(r,\delta)=(j+1,n_{2}-j)$-LRC, $Q^{\prime}(C_{\Delta}^{P})$.
The parameters of $Q^{\prime}(C_{\Delta}^{P})$ are $[[n_{1}n_{2},2[(j+1)(n_{1}-1)+s+1]-n_{1}n_{2},\geq(n_{2}-s)]]_{q}$.
[…]
(a) Sets $\Delta_{i,s}^{2}$
[…]
(b) Sets $\Delta_{j,s}^{2,\sigma}$
Figure 2. Sets $\Delta_{i,s}^{2}$ and $\Delta_{j,s}^{2,\sigma}$ in Proposition 33
Example 2.
Consider the values $q=7$, $n_{1}=n_{2}=7$, $i=5$ and $j=6$. Set $\Delta=\Delta_{5,6}$ as defined in Proposition 32. Since $i=5\geq\frac{6}{2}$, Proposition 32 proves that $Q^{\prime}(C_{\Delta}^{P})$ is an optimal quantum $(r,\delta)=(6,2)$-LRC with parameters $[[49,35,\geq 2]]_{7}$. Notice that $Q^{\prime}(C_{\Delta}^{P})$ is also a quantum LRC with locality $6$ attaining the bound (1).
5.2. Some examples of optimal quantum $(r,\delta)$-LRCs by using Hermitian duality
We are going to show that $(r,\delta)$-LRCs coming from Hermitian dual-containing MDS $q^{2}$-ary classical codes provide optimal quantum $(r,\delta)$-LRCs. Note that the same happens when considering MDS $q$-ary classical codes and Euclidean dual containment.
Proposition 34.
Let $C$ be an $[n,k,n-k+1]_{q^{2}}$ Hermitian dual-containing classical code. Then, $C$ is an optimal $(r,\delta)=(k,n-k+1)$-LRC that satisfies $\delta\leq d(C^{\perp_{h}})$, Therefore, by Theorem 28, $Q^{\prime}(C)$ is an optimal quantum $(r,\delta)=(k,n-k+1)$-LRC with parameters $[[n,2k-n,n-k+1]]_{q}$.
Proof.
$C$ is an MDS code and thus $C^{\perp_{h}}$ has parameters $[n,n-k,k+1]_{q^{2}}$. In addition, $C$ is an $(r,\delta)=(k,n-k+1)$-LRC because $r+\delta-1=n$ and the parity-check matrix of $C$ has size $(n-k)\times n$ and maximum rank, which shows that it allows to recover $\delta-1$ erasures. Now, $C^{\perp_{h}}\subseteq C$ proves inequality $k\geq\frac{n}{2}$ and then $\delta=n-k+1\leq d(C^{\perp_{h}})=k+1$. The fact that $Q^{\prime}(C)$ is an optimal quantum $(r,\delta)=(k,n-k+1)$-LRC follows from Theorems 28 and 30. Finally, we get the exact distance of the quantum code from the quantum Singleton bound. This concludes the proof. ∎
Remark 35.
Proposition 34 (and its obvious analogue for MDS $q$-ary classical codes) show that any MDS quantum stabilizer code, obtained from a classical code satisfying Hermitian (or Euclidean) dual containment, is optimal for the locality described in the statement. This is because these stabilizer codes always come from a dual-containing code where $2k-n\geq 0$. Specific parameters of many MDS stabilizer codes can be consulted in [54] and references therein.
Conflict of interest
The authors declare they have no conflict of interest.